# Preprint submitted to Journal

### v4

@@ 3.1.2 Properties
Definition 8 (The bound 2-tuple of a non-empty $\mathrm{2Tuple}\mathcal{T}$ instance with limited length & helper function & permutation invariance).
@@ 3.2 Set of optional intervals events
Definition 12 (Set of feasible interval 2-tuple sets of finite $\mathrm{OIE}\mathcal{S}$ instance).
@@
$\{\mathcal{F}_{\mathrm{oie}_{1}},\mathcal{F}_{\mathrm{oie}_{2}},\cdots,\mathcal{F}_{\mathrm{oie}_{n}}\}$
is called “The set of feasible interval 2-tuple sets of $\mathrm{oie}\mathcal{S}$”, abbreviated as $\mathcal{FS}$ of $\mathrm{oie}\mathcal{S}$.
For the special case where an $\mathrm{OIE}\mathcal{S}$ instance is $\emptyset$, its $\mathcal{FS}$ is $\emptyset$.
For non-empty $\mathrm{OIE}\mathcal{S}$ instances, $\mathcal{FS}$ is an instance of the $\mathrm{2Tuple}\mathcal{SS}$ type mentioned earlier.
@@
Then, its $\mathcal{FS}$ is
$\mathcal{FS}_{\mathrm{oie}\mathcal{S}_{Drs}}$
$\quad=\{\ \mathcal{F}_{\mathrm{oie}_{Dr_{A}}},\ \mathcal{F}_{\mathrm{oie}_{Dr_{B}}},\ \mathcal{F}_{\mathrm{oie}_{Dr_{C}}}\ \}$
@@ 4 Two sequential operations on the elements of a finite set of optional intervals events
We first carry out the work of using the set indices to represent the operation order of the elements of the set.
@@
Then the tuple $(i_{1},i_{2},i_{3},\cdots,i_{n})$ is an index tuple of the finite index set $\mathrm{I}_{\mathcal{S}}$, denoted as $\mathrm{Idx}\mathcal{T}$ with length nn”.
@@
Now, we need to design a function to get a set that is naturally isomorphic to the Cartesian product of all the elements in a finite $\mathrm{2Tuple}\mathcal{SS}$(set of sets of 2-tuples) instance, and this function serves as a helper function for refining the sequential operation rules of $\mathrm{OIE}\mathcal{S}$ later.
@@
What we want is not the result restricted by parentheses in terms of operation order, but the flattened result of each element in this resulting set. This is why we use natural isomorphism. The significance of natural isomorphism will become more apparent later when analyzing the associativity of temporal operations.
@@
$\mathrm{2tuple}\mathcal{TS}^{NatIso2CP}=\mathfrak{f}\mathrm{NatIso2CP}(\mathcal{FS}_{\mathrm{oie}\mathcal{S}},\mathrm{idx}\mathcal{T})$
@@
The purpose of calculating set that is naturally isomorphic to the Cartesian product of all third items($\mathfrak{m}\mathrm{Int2tuple}\mathcal{S}$) of the members of an $\mathrm{OIE}\mathcal{S}$ instance, is to calculate the maximum combination as for subsequent use. However, since $\mathrm{OIE}$ instances are used to map Event instances, and Event instances are realized in our real world, which is a world with various limitations. As a result, for this $\mathrm{2Tuple}\mathcal{TS}$ instance, we also need to discuss some limiting factors. The content to be discussed in this subsection is that, considering this result as a set, we need to identify which elements of the set are infeasible, and then exclude them. The remaining elements of the set are the available ones, and they also have their subsequent use in operations.
@@
For an $\mathrm{OIE}$ instance $\mathrm{oie}$, when analyzing its execution either independently or in combination with other $\mathrm{OIE}$ instances, Whether each interval in the third item of $\mathrm{oie}$ can be implemented may change. That is, because there is independence and dependence between Event instances, $\mathrm{OIE}$ instances will also have the same properties.
Given that $\mathrm{oie}$ already has its own 3rd element $\mathcal{I}$, once the existence of other $\mathrm{OIE}$ instances is introduced, The feasibility of each interval in $\mathfrak{m}\mathrm{Int2tuple}\mathcal{S}_{\mathrm{oie}}$ may vary, that is, some intervals can no longer be implemented in reality.
@@
$\mathrm{2tuple}\mathcal{TS}^{NatIso2CP}=\mathfrak{f}\mathrm{NatIso2CP}(\mathcal{FS}_{\mathrm{oie}\mathcal{S}_{father-son}},(\mathrm{idx}_{father},\mathrm{idx}_{son}))$
@@
$\mathrm{2tuple}\mathcal{TS}^{NatIso2CP}=\mathfrak{f}\mathrm{NatIso2CP}(\mathcal{FS}_{\{\mathrm{oie}\mathcal{S}\}},\mathrm{idx}\mathcal{T})$
@@
Now, we have already used Definition 15 to calculate the set that is naturally isomorphic to the Cartesian product for all members of the $\mathcal{FS}$ instance of the $\mathrm{OIE}\mathcal{S}$ instance, and Definition 16 has provided the set to which the infeasible elements belong. So we can narrow down the scope, we only need to remove the infeasible elements from set that is naturally isomorphic to the Cartesian product to get the usable elements.
@@
. Take the set that is naturally isomorphic to the Cartesian product of all members of the $\mathcal{FS}$ of $\mathrm{oie}\mathcal{S}$ in the index order specified by $\mathrm{idx}\mathcal{T}$
$\mathrm{2tuple}\mathcal{TS}^{NatIso2CP}=\mathfrak{f}\mathrm{NatIso2CP}(\mathcal{FS}_{\mathrm{oie}\mathcal{S}},\mathrm{idx}\mathcal{T})$
@@ 4.3 Complete sequential addition
$\mathrm{domFlt2tuple}=(\mathrm{ts}_{\theta},\mathrm{te}_{\theta})$
@@
$\mathcal{C}_{\mathrm{oie}_{result}}$
$\qquad=flatten(oie_{\mathrm{idx}_{1}},\ \mathrm{oie}_{\mathrm{idx}_{2}},\ \mathrm{oie}_{\mathrm{idx}_{3}},\ \cdots,\ \mathrm{oie}_{\mathrm{idx}_{n}})$
$\qquad=(\mathrm{oie}_{\mathrm{idx}_{1}},\ \mathrm{oie}_{\mathrm{idx}_{2}},\ \mathrm{oie}_{\mathrm{idx}_{3}},\ \cdots,\ \mathrm{oie}_{\mathrm{idx}_{n}})$
@@
Step 3. Using Definition 19, get the corresponding $\mathrm{2Tuple}\mathcal{TS}_{\textbf{Dom}}\big|_{\mathrm{ts}_{\theta}}^{\mathrm{te}_{\theta}}$ instance
$\mathrm{2tuple}\mathcal{TS}_{dom}\big|_{\mathrm{ts}_{\theta}}^{\mathrm{te}_{\theta}}$ $=\mathfrak{f}\mathrm{DomFlt}\mathrm{2tuple}\mathcal{TS}(\mathrm{2tuple}\mathcal{TS}^{feas},\mathrm{ts}_{\theta},\mathrm{te}_{\theta})$
$=\mathfrak{f}\mathrm{DomFlt}\mathrm{2tuple}\mathcal{TS}(\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}),\mathrm{ts}_{\theta},\mathrm{te}_{\theta})$
@@
$\mathrm{2tuple}\mathcal{TS}_{dom}\big|_{\mathrm{ts}_{\theta}}^{\mathrm{te}_{\theta}}=\{\mathrm{2tuple}\mathcal{T}_{1},\mathrm{2tuple}\mathcal{T}_{2},\cdots,\mathrm{2tuple}\mathcal{T}_{\lambda}\}$
@@
$\mathrm{2tuple}\mathcal{TS}_{dom}\big|_{\mathrm{ts}_{\theta}}^{\mathrm{te}_{\theta}}\neq\emptyset$
then the $\mathcal{F}$(2nd element) of $\mathrm{oie}_{result}$ is assigned to $\mathrm{2tuple}\mathcal{TS}_{dom}\big|_{\mathrm{ts}_{\theta}}^{\mathrm{te}_{\theta}}$
$\mathcal{F}_{\mathrm{oie}_{result}}=\mathrm{2tuple}\mathcal{TS}_{dom}\big|_{\mathrm{ts}_{\theta}}^{\mathrm{te}_{\theta}}$
@@
$\mathrm{2tuple}\mathcal{TS}_{dom}\big|_{\mathrm{ts}_{\theta}}^{\mathrm{te}_{\theta}}=\emptyset$
@@
$\mathcal{I}_{\mathrm{oie}_{result}}$ $=\bigg\{\ \mathfrak{f}\mathrm{Bound2tuple}(\mathrm{2tuple}\mathcal{T}_{i})\ \bigg|\mathrm{2tuple}\mathcal{T}_{i}\in\mathrm{2tuple}\mathcal{TS}_{dom}\big|_{\mathrm{ts}_{\theta}}^{\mathrm{te}_{\theta}}\ \bigg\}$
@@
$\qquad\quad\mathrm{2tuple}\mathcal{T}_{i}\in\mathfrak{f}\mathrm{DomFlt}\mathrm{2tuple}\mathcal{TS}(\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}),\mathrm{ts}_{\theta},\mathrm{te}_{\theta})\ \bigg\}$
@@
$\qquad\mathfrak{f}\mathrm{DomFlt}\mathrm{2tuple}\mathcal{TS}(\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}),\mathrm{ts}_{\theta},\mathrm{te}_{\theta}),$
@@
$\qquad\qquad\mathrm{2tuple}\mathcal{T}_{i}\in\mathfrak{f}\mathrm{DomFlt}\mathrm{2tuple}\mathcal{TS}(\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}),\mathrm{ts}_{\theta},\mathrm{te}_{\theta})\ \bigg\},$
@@
After executing the above-mentioned process, including the process where the result is $\mathrm{oie}_{\mathrm{void}}$, it is called “The complete sequential addition of members of a finite $\mathrm{OIE}\mathcal{S}$ instance under domain-filtering 2-tuple $(\mathrm{ts}_{\theta},\mathrm{te}_{\theta})$ in index order $\mathrm{idx}\mathcal{T}$”, denoted as

### v5

@@ 3.1.2 Properties
Definition 7 (The bound 2-tuple of a finite non-empty $\mathrm{2Tuple}\mathcal{T}$ instance & helper function & permutation invariance).
@@ 3.2 Set of optional intervals events
Definition 11 (Set of $\mathcal{I}$ of finite $\mathrm{OIE}\mathcal{S}$ instance).
@@
$\{\mathcal{I}_{\mathrm{oie}_{1}},\mathcal{I}_{\mathrm{oie}_{2}},\cdots,\mathcal{I}_{\mathrm{oie}_{n}}\}$
is called “The set of $\mathcal{I}$ of $\mathrm{oie}\mathcal{S}$”, abbreviated as $\mathcal{IS}$ of $\mathrm{oie}\mathcal{S}$.
For the special case where an $\mathrm{OIE}\mathcal{S}$ instance is $\emptyset$, its $\mathcal{IS}$ is $\emptyset$.
For non-empty $\mathrm{OIE}\mathcal{S}$ instances, $\mathcal{IS}$ is an instance of the $\mathrm{2Tuple}\mathcal{SS}$ (set of sets of 2-tuples) type.
@@
Then, its $\mathcal{IS}$ is
$\mathcal{IS}_{\mathrm{oie}\mathcal{S}_{Drs}}$
$\quad=\{\ \mathcal{I}_{\mathrm{oie}_{Dr_{A}}},\ \mathcal{I}_{\mathrm{oie}_{Dr_{B}}},\ \mathcal{I}_{\mathrm{oie}_{Dr_{C}}}\ \}$
@@ 4 Two sequential operations on the elements of a finite set of optional intervals events
We first carry out the work of using the set indices to represent the operation order of the elements of a set.
@@
Then the tuple $(i_{1},i_{2},i_{3},\cdots,i_{n})$ is an index tuple of the finite index set $\mathrm{I}_{\mathcal{S}}$, denoted as $\mathrm{Idx}\mathcal{T}$” with length $n$.
@@
Now, we need to design a function to get a set that is naturally isomorphic to the Cartesian product of all the elements in a finite $\mathrm{2Tuple}\mathcal{SS}$(set of sets of 2-tuples) instance, and this function serves as a helper function for refining the sequential operation rules later.
@@
What we want is not the result restricted by parentheses in terms of operation order, but the flattened result of each element in this resulting set. This is why we use natural isomorphism.
@@
$\mathrm{2tuple}\mathcal{TS}^{NatIso2CP}=\mathfrak{f}\mathrm{NatIso2CP}(\mathcal{IS}_{\mathrm{oie}\mathcal{S}},\ \mathrm{idx}\mathcal{T})$
@@
The purpose of calculating set that is naturally isomorphic to the Cartesian product of all $\mathcal{I}$(3rd elements) of the members of an $\mathrm{OIE}\mathcal{S}$ instance, is to calculate the maximum combination as for subsequent use. However, since $\mathrm{OIE}$ instances are used to map Event instances, and Event instances are realized in our real world, which is a world with various limitations. As a result, for this $\mathrm{2Tuple}\mathcal{TS}$ instance, we also need to discuss some limiting factors. The content to be discussed in this subsection is that, considering this result as a set, we need to identify which elements of the set are infeasible, and then exclude them. The remaining elements of the set are the available ones.
@@
For an $\mathrm{OIE}$ instance $\mathrm{oie}$, when analyzing its execution either independently or in combination with other $\mathrm{OIE}$ instances, Whether each interval in the $\mathcal{I}$(3rd element) of $\mathrm{oie}$ can be implemented may change. That is, because there is independence and dependence between Event instances, $\mathrm{OIE}$ instances will also have the same properties.
Given that $\mathrm{oie}$ already has its own $\mathcal{I}$(3rd element), once the existence of other $\mathrm{OIE}$ instances is introduced, The feasibility of each interval in $\mathcal{I}_{\mathrm{oie}}$ may vary, that is, some intervals can no longer be implemented in reality.
@@
$\mathrm{2tuple}\mathcal{TS}^{NatIso2CP}=\mathfrak{f}\mathrm{NatIso2CP}(\mathcal{IS}_{\mathrm{oie}\mathcal{S}_{father-son}},\ (\mathrm{idx}_{father},\mathrm{idx}_{son}))$
@@
$\mathrm{2tuple}\mathcal{TS}^{NatIso2CP}=\mathfrak{f}\mathrm{NatIso2CP}(\mathcal{IS}_{\{\mathrm{oie}\mathcal{S}\}},\ \mathrm{idx}\mathcal{T})$
@@
Now, we have already used Definition 14 to calculate the set that is naturally isomorphic to the Cartesian product for all members of the $\mathcal{IS}$ instance of the $\mathrm{OIE}\mathcal{S}$ instance, and Definition 15 has provided the set to which the infeasible elements belong. So we can narrow down the scope, we only need to remove the infeasible elements from set that is naturally isomorphic to the Cartesian product to get the usable elements.
@@
. Take the set that is naturally isomorphic to the Cartesian product of all members of the $\mathcal{IS}$ of $\mathrm{oie}\mathcal{S}$ in the index order specified by $\mathrm{idx}\mathcal{T}$
$\mathrm{2tuple}\mathcal{TS}^{NatIso2CP}=\mathfrak{f}\mathrm{NatIso2CP}(\mathcal{IS}_{\mathrm{oie}\mathcal{S}},\ \mathrm{idx}\mathcal{T})$
@@
4.2.3 Mutually independent $\mathrm{OIE}\mathcal{S}$
Definition 17 (Mutually independent $\mathrm{OIE}\mathcal{S}$).
Suppose there is an $\mathrm{OIE}\mathcal{S}$ instance $\mathrm{oie}\mathcal{S}$ with cardinality $n\in\mathbb{N}^{+}$, with form
and every element therein must not be $\mathrm{oie}_{\mathrm{void}}$. Let
$\mathrm{idx}\mathcal{T}_{asc}=(1,2,3,\cdots,n)$
If
$\mathfrak{f}\mathrm{Inf2tupleTS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}_{asc})=\emptyset$
then $\mathrm{oie}\mathcal{S}$ is called mutually independent $\mathrm{oie}\mathcal{S}$. Such instances are called “Mutually Independent $\mathrm{OIE}\mathcal{S}$”, denoted as $\mathrm{OIE}\mathcal{S}^{\text{indep}}$.
Scenarios including 3 doctors submitting their manuscripts via their respective computers, multiple computing nodes in a supercomputing center executing tasks without mutual interference, and several athletes competing in their individual lanes without affecting one another, can all be characterized by Definition 17. And many subsequent examples in this paper will also be based on Definition 17.
@@ 4.3 Complete sequential addition
$\mathrm{domFlt2tuple}=(\alpha,\beta)$
@@
Step 3. Using Definition 19, get the corresponding $\mathrm{2Tuple}\mathcal{TS}_{\textbf{Dom}}\big|_{\alpha}^{\beta}$ instance
$\mathrm{2tuple}\mathcal{TS}_{dom}\big|_{\alpha}^{\beta}$ $=\mathfrak{f}\mathrm{DomFlt}\mathrm{2tuple}\mathcal{TS}(\mathrm{2tuple}\mathcal{TS}^{feas},\alpha,\beta)$
$=\mathfrak{f}\mathrm{DomFlt}\mathrm{2tuple}\mathcal{TS}(\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}),\alpha,\beta)$
@@
$\mathrm{2tuple}\mathcal{TS}_{dom}\big|_{\alpha}^{\beta}=\{\mathrm{2tuple}\mathcal{T}_{1},\mathrm{2tuple}\mathcal{T}_{2},\cdots,\mathrm{2tuple}\mathcal{T}_{\lambda}\}$
@@
$\mathrm{2tuple}\mathcal{TS}_{dom}\big|_{\alpha}^{\beta}\neq\emptyset$
then the $\mathcal{F}$(2nd element) of $\mathrm{oie}_{result}$ is assigned to $\mathrm{2tuple}\mathcal{TS}_{dom}\big|_{\alpha}^{\beta}$
$\mathcal{F}_{\mathrm{oie}_{result}}=\mathrm{2tuple}\mathcal{TS}_{dom}\big|_{\alpha}^{\beta}$
@@
$\mathrm{2tuple}\mathcal{TS}_{dom}\big|_{\alpha}^{\beta}=\emptyset$
@@
$\mathcal{I}_{\mathrm{oie}_{result}}$ $=\bigg\{\ \mathfrak{f}\mathrm{Bound2tuple}(\mathrm{2tuple}\mathcal{T}_{i})\ \bigg|\mathrm{2tuple}\mathcal{T}_{i}\in\mathrm{2tuple}\mathcal{TS}_{dom}\big|_{\alpha}^{\beta}\ \bigg\}$
@@
$\qquad\quad\mathrm{2tuple}\mathcal{T}_{i}\in\mathfrak{f}\mathrm{DomFlt}\mathrm{2tuple}\mathcal{TS}(\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}),\alpha,\beta)\ \bigg\}$
@@
$\qquad\mathfrak{f}\mathrm{DomFlt}\mathrm{2tuple}\mathcal{TS}(\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}),\alpha,\beta),$
@@
$\qquad\qquad\mathrm{2tuple}\mathcal{T}_{i}\in\mathfrak{f}\mathrm{DomFlt}\mathrm{2tuple}\mathcal{TS}(\mathfrak{f}\mathrm{Feas2tuple}\mathcal{TS}(\mathrm{oie}\mathcal{S},\mathrm{idx}\mathcal{T}),\alpha,\beta)\ \bigg\},$
@@
After executing the above-mentioned process, including the process where the result is $\mathrm{oie}_{\mathrm{void}}$, it is called “The complete sequential addition of members of a finite $\mathrm{OIE}\mathcal{S}$ instance under domain-filtering 2-tuple $(\alpha,\beta)$ in index order $\mathrm{idx}\mathcal{T}$”, denoted as
@@ 4.6.3 Permutationally equivalent relation of results of complete sequential multiplication
4.7 Natural Complete sequential addition
Definition 24 (Natural complete sequential addition of members of a finite $\mathrm{OIE}\mathcal{S}$ instance in an index order).
There is a finite $\mathrm{OIE}\mathcal{S}$ instance $\mathrm{oie}\mathcal{S}$ containing $n(n\in\mathbb{N}^{+})$ elements in the ordered enumeration form
and an $\mathrm{Idx}\mathcal{T}$ instance with length $n$
For any $\mathrm{OIE}\mathcal{S}$ instance $\mathrm{oie}_{\lambda}$, let
$\pi_{1}^{\min}(\mathrm{oie}_{\lambda})=\min\left\{\pi_{1}(t)\mid t\in\mathcal{I}_{\mathrm{oie}_{\lambda}}\right\}$
where $\pi_{1}(t)$ denotes the first projection(get the 1st element) of 2-tuple $t$, and
$\pi_{2}^{\max}(\mathrm{oie}_{\lambda})=\max\left\{\pi_{2}(t)\mid t\in\mathcal{I}_{\mathrm{oie}_{\lambda}}\right\}$
where $\pi_{2}(t)$ denotes the second projection(get the 2nd element) of 2-tuple $t$.
If
$\exists(A,B),\forall i\in[1,n]:$
$\qquad\mathrm{oie}_{i}\neq\mathrm{oie}_{\mathrm{void}}\land\pi_{1}^{\min}(\mathrm{oie}_{i})=A\land\pi_{2}^{\max}(\mathrm{oie}_{i})=B$
then by Definition 20, the following complete sequential addition
${\oplus|_{A}^{B}}(\mathrm{oie}_{\mathrm{idx}_{1}},\ \mathrm{oie}_{\mathrm{idx}_{2}},\ \cdots,\ \mathrm{oie}_{\mathrm{idx}_{n}})$
is called “Natural Complete Sequential Addition”, and the expression can be simplified to
$\oplus(\mathrm{oie}_{\mathrm{idx}_{1}},\ \mathrm{oie}_{\mathrm{idx}_{2}},\ \cdots,\ \mathrm{oie}_{\mathrm{idx}_{n}})$
Natural complete sequential addition serves as a natural special case of complete sequential addition free from subjective intervention, which eliminates the subjectivity arising from the artificial specification of the domain-filtering 2-tuple. It directly takes the common and natural time boundaries of all $\mathrm{OIE}$ instances participating in the operation, namely the unified minimum left boundary A and the unified maximum right boundary B, as the filtering domain of the operation, thus ensuring the objectivity and natural adaptability of the operation.
It fully inherits the core semantics of complete sequential addition and is still applied to model concurrent events with equal opportunity within a shared time domain. It only imposes a naturalized and standardized qualification on the filtering domain without altering the core rules of the operation. Meanwhile, it provides a concise form for the standardized modeling of real-world scenarios, fitting well with concurrent scenarios that naturally share a unified time window such as the 100 metres race. For such scenarios, event modeling can be completed directly through the simplified expression without additional specification of the filtering domain.
From a formal perspective, the removal of the filtering domain in Natural Complete Sequential Addition results in a notational form identical to that of Complete Sequential Multiplication. This unification endows the two sequential operations with a high degree of mathematical elegance.